OpenStudy (anonymous):

\frac{2 x}{x^2-4}=\frac{4}{x^2-4}-\frac{3}{x+2 } Will someone solve this?

OpenStudy (anonymous):

\[{2x \over x^2 - 4} = {4 \over x^2 - 4} - {3 \over x + 2}\] ?

OpenStudy (anonymous):

\[{2x \over (x+2)(x-2)} = {4 \over (x+2)(x-2)} - {3 \over x + 2}\]

OpenStudy (anonymous):

Multiply each side by (x + 2)(x - 2) to get: \[2x = 4 - 3(x-2)\] solve from there